OpenStudy (anonymous):

solve for x: 5^(x+1) = 6^(2x) I think I have to get the bases the same, but I don't know how.

OpenStudy (anonymous):

Take log to the both sides .....

OpenStudy (anonymous):

I don't get it can you explain @Eyad

OpenStudy (anonymous):

when u take log for the both sides ,the power will get down : (x+1) log5=(2x) log6 (x+1)/(2x)=Log6/log5...... Get it